# OATS: Outlier-Aware Pruning through Sparse and Low Rank Decomposition

### v1

@@ 3.3 Ablations for OATS
The results of our ablations are depicted in Table 6 below:
@@
Table 6: Ablations of OATS on Phi-3-Mini at $40\%$ compression rate with a rank ratio of $20\%$. The experiments show that scaling the weights by the second moment of the input activations and pruning row-wise significantly improves the performance of OATS.
In addition to the ablations, in Figure 1 below, we show how the rank ratio and the number of iterations impact the performance of OATS .
0.10.20.30.40.50.60.74040404050505050606060607070707080808080Rank RatioAccuracy (%)120406080100120140160Number of Iterations0-shotMMLU
@@
The experiments reveal that the performance under the rank ratio generally saturates between $25\%$ to $30\%$ prior to deteriorating. For the number of iterations, there is a sharp increase in performance from 00-$20$ iterations before leveling off and saturating at around $80$ iterations.
@@ 3.4 Hardware Speedup
We utilize the DeepSparse engine by NeuralMagic (2021) to measure the CPU acceleration induced by OATS compared to pruned models with unstructured sparsity. We run end-to-end inference on a compressed Phi-3 Medium 15B model for a single batch of 2048 tokens on an Intel Xeon Gold 6148 CPU @ 2.40GHz with 32 cores. The achieved speed-ups are depicted in Table 7 below.
Compression Rate Method Throughput (tokens/sec) Speedup
Dense 4.03 1.00×1.$0$
@@
Table 7: Comparison of throughput and speedup achieved through OATS and unstructured pruning methods relative to their dense counterparts.
Our results highlight that by trading unstructured sparsity for structured sparsity through the low-rank terms, OATS leads to significantly more CPU speed-up compared to models that rely solely on unstructured pruning. In particular, at $40\%$ compressed, OATS has 1.37×1.37$\times$1.37 × the speed-up than unstructured pruning.
@@
We benchmark the N:M performance of OATS against the 2:4 structured sparsity performance of current state of the art pruning algorithms. Since the low-rank terms found by OATS remains dense, we opt to impose the sparser structured sparsity pattern of 2:8 on the sparse term returned by OATS testing various rank ratios of $0.25$, $0.3$, $0.35$, $0.4$, $0.45$, and $0.5$. In Figure 2 below, we depict the compression ratio versus the 5-shot MMLU accuracy of different compression algorithms.
5054586164663030303040404040MMLU Avg. AccuracyPhi-3 Mini505458616466Llama-3 8BOATSSparseGPTWandaDSNoT
@@
Our results show that despite having a much stricter structured sparsity pattern of 2:8, OATS is able to recover the model performance through the presence of its low-rank terms. Specifically, at a fixed compression rate of $50\%$, OATS is able to outperform all prior state-of-the-art by up to $9.2\%$. In the case of Llama-3 8B, OATS surpasses previous methods by $2.86\%$ with $8\%$ more compression.
@@ 4 Experiments on Vision Transformers
We run experiments on Google’s ViT-Base (Wu et al., 2020), an $86.6$M parameter model trained in a supervised manner on ImageNet-21k (Ridnik et al., 2021) and fine-tuned on ImageNet 2012 (Russakovsky et al., 2015), and DinoV2-Giant (Oquab et al., 2023), a $1.14$B parameter model that was trained utilizing self-supervised learning.
We benchmark OATS against the same three pruning algorithms: SparseGPT, Wanda, and DSNoT, by evaluating on the validation split of ImageNet (Russakovsky et al., 2015). A subset of size $2048$ from the training split of ImageNet is utilized for the calibration data and is maintained consistent across all pruning experiments. All OATS experiments utilize a rank ratio of $\kappa=20\%$ and $N=80$ iterations. We exclude from compression the embedding layers and the classifier layer.
Depicted in Table 8 below are the validation accuracies of ImageNet for ViT-Base and DinoV2-Giant under various compression rates.
Compression Rate Method ViT-Base DinoV2-Giant
@@ 5 Visualizing and Interpreting the Decomposition
To develop a better understanding of how the sparse and low rank components individually contribute to the flow of information through the attention mechanism, we compute and visualize the attention rollout (Abnar & Zuidema, 2020) of the compressed model when:
@@
Figure 3: A visualization of how the attention rollout is computed to isolate the contribution of the sparse terms versus low-rank terms given by the OATS compression algorithm.
We utilize various images in the Microsoft COCO dataset (Lin et al., 2014) to generate the attention rollout. Figure 4 below depicts the attention rollout computed using the scheme described above on ViT-B that was compressed by $50\%$ with a rank ratio of $20\%$.
@@
The rollout visualizations show that the sparse and low-rank terms capture distinct areas of the image, effectively segmenting it, and reveal three distinct patterns in how the image is partitioned. The first type of behavior, which is also commonly exhibited in the classical setting (Candès et al., 2011), is when one component (commonly sparse) captures the subject(s), while the other component (commonly low-rank) captures the background. The second behavior occurs when both components focus on different parts of the same subject, capturing distinct features. The third behavior arises when the image contains multiple subjects, with each component isolating a different subject. While these patterns provide initial insights into how the components process visual information, further investigation is needed to fully understand the underlying mechanisms driving these behaviors.
@@ 6 Related Works
OATS utilizes the same outlier scaling as the the one that is employed by Wanda (Sun et al., 2024b). In fact, Wanda can be seen as a special case of OATS when the rank ratio $\kappa=0$, which, according to Equation 2, results in the low-rank term becoming the $\mathbf{0}$ matrix. This leads to OATS performing a singular hard thresholding step that is equivalent to the pruning step described by Wanda and returning a compressed matrix equal to:
${\bm{W}}_{\text{compressed}}=\textsc{HardThreshold}({\bm{W}}{\bm{D}},k){\bm{D}}^{-1}.$
@@
The emergence of sparse and low-rank structures in transformers has recently become an area of both theoretical and practical interest. On the theoretical front, Zhao et al. (2024c) showed that the logits of LLMs trained utilizing next token prediction converge to a low rank and sparse structure. On the practical front, Scatterbrain proposed by Chen et al. (2021) shows that its possible to approximate the entire attention mechanism with a single sparse and low rank representation. LoRAP by Li et al. (2024) finds a low-rank representation for the attention matrices by utilizing a similar scaling technique as OATS. Recent works, such as LoSparse (Li et al., 2023), LoRAPrune (Zhang et al., 2024a), and APT (Zhao et al., 2024a), propose variations of applying structured pruning on the weights while incorporating a low-rank adapter that is trained via gradient descent. We want to highlight that unlike OATS, these works utilize training and low-rank adaptation to recover model performance from structured pruning, whereas OATS depends on an intrinsic sparse and low-rank approximation of the weight matrices.
@@
While OATS utilizes an alternating thresholding approach for its efficiency and simplicity, the search for fast Robust PCA algorithms has been a key area of interest since the inception of the problem. An alternative approach is to apply a convex relaxation, replacing the sparsity and low-rank constraints with an $\ell_{1}$ and nuclear norm surrogate leading to Stable Principle Component Pursuit (Zhou et al., 2010). Another way to enhance computational efficiency is to utilize gradient descent by parameterizing the low-rank matrix as ${\bm{L}}={\bm{U}}{\bm{V}}^{\top}$, and then performing gradient descent on ${\bm{U}}$ and ${\bm{V}}$ (Yi et al., 2016; Tong et al., 2021).
@@
While pruning has been proven to be a successful approach to model compression, there is still ongoing research into what pruning is pruning and how it impacts model performance. Paganini (2020) show that magnitude-based pruning has a disproportionate negative effect on underrepresented classes. In a similar vein, Yin et al. (2024a) show that pruning LLMs can lead to irreversible damage to model performance on tasks that are more challenging. We highlight that the low-rank term present in OATS might be able to mitigate the negative impacts of pruning raised by the prior works. Indeed, in Tables 2 and 5, we can see that at higher compression rates, the gap between OATS and prior state-of-the-art is significantly larger suggesting that the low-rank term plays a critical role in mitigating the loss in performance at high compression.
@@ 7 Conclusion
We have introduced OATS, an algorithm that distinguishes itself from prior algorithms by utilizing outlier information to determine a compressed sparse and low-rank decomposition without any re-training. We run a number of experiments showing that OATS is able to consistently outperform prior state-of-the-art across multiple benchmarks and compression rates while also improving on speed-up. Furthermore, in the search for interpretability, we have generated visualizations highlighting that the sparse and low-rank terms capture different segments of the image. We hope that our work can be utilized to investigate how sparse and low-rank structures in transformers can be fully leveraged for maximal memory and compute efficiency.
@@ Appendix A Appendix
A.1 Zero-Shot Task-Specific Performance
Table 9 depicts the task-specific performance for the zero-shot evaluation results presented in Section 3.2 and Appendix A.4.
Model Compression Rate Method PIQA HellaSwag WinoGrande OpenBookQA RTE BoolQ ARC-e ARC-c Phi-3 Mini 0% Dense 81.23 77.50 73.56 46.80 75.81 85.32 78.45 57.25 30% SparseGPT 78.94 76.94 69.85 49.60 73.29 84.13 76.39 55.89 Wanda 79.65 76.27 71.59 48.00 73.65 83.70 77.23 55.20 DSNoT w/ SparseGPT 80.09 75.61 72.22 47.40 74.37 84.22 77.86 54.69 DSNoT w/ Wanda 80.41 75.52 72.06 47.60 74.37 84.53 79.55 55.55 OATS 80.03 77.07 72.61 47.60 74.37 84.92 77.44 57.76 40% SparseGPT 78.35 75.07 68.59 47.00 72.20 83.67 75.29 53.24 Wanda 78.35 73.87 69.30 45.40 71.84 83.18 76.52 51.96 DSNoT w/ SparseGPT 78.56 72.99 70.64 46.20 70.40 82.72 76.52 52.82 DSNoT w/ Wanda 79.33 73.06 70.88 44.00 70.76 83.79 77.40 53.41 OATS 79.38 75.86 70.01 46.60 72.56 83.98 76.85 55.12 50% SparseGPT 77.20 70.63 66.46 45.20 70.76 83.06 70.58 47.01 Wanda 76.33 67.70 66.38 41.80 66.43 81.83 72.43 47.35 DSNoT w/ SparseGPT 76.28 67.16 65.90 42.20 63.90 81.56 72.90 48.04 DSNoT w/ Wanda 75.52 66.54 67.64 43.00 65.34 82.54 73.48 48.55 OATS 77.26 71.64 69.53 44.80 73.65 81.28 77.10 52.05 Phi-3 Medium 0% Dense 81.66 82.83 75.85 50.00 77.62 88.17 78.41 59.64 30% SparseGPT 81.39 82.02 75.77 50.80 77.26 87.80 80.05 61.18 Wanda 81.39 80.88 76.01 49.40 76.90 87.74 79.59 60.49 DSNoT w/ SparseGPT 81.94 80.76 76.95 48.40 75.81 87.65 79.25 59.81 DSNoT w/ Wanda 81.66 81.03 77.27 49.20 76.53 87.80 78.96 59.81 OATS 81.07 82.09 74.43 51.20 78.34 88.38 78.16 58.70 40% SparseGPT 80.41 80.70 75.53 51.20 77.26 88.32 81.23 60.58 Wanda 79.87 78.15 75.45 48.60 77.26 87.71 78.11 58.96 DSNoT w/ SparseGPT 79.82 78.07 75.37 47.00 76.53 87.98 77.31 58.19 DSNoT w/ Wanda 80.30 78.11 74.66 47.80 77.26 88.04 78.11 58.87 OATS 81.39 81.72 75.06 51.00 77.62 87.65 80.39 60.84 50% SparseGPT 79.71 78.27 73.64 50.40 75.45 87.09 82.03 59.39 Wanda 78.29 74.07 74.03 45.00 75.81 85.72 77.44 57.34 DSNoT w/ SparseGPT 79.27 74.30 74.59 44.40 76.90 85.26 77.69 56.57 DSNoT w/ Wanda 78.56 73.81 75.14 43.60 75.81 86.33 77.53 58.02 OATS 81.07 79.18 76.09 50.20 74.73 87.77 80.05 58.02 Llama-3 8B 0% Dense 80.74 79.16 73.40 45.0 0 67.87 80.98 77.69 53.50 30% SparseGPT 80.36 78.58 73.24 44.40 66.79 81.38 76.81 51.11 Wanda 79.98 78.00 73.64 44.40 64.26 81.62 76.18 50.94 DSNoT w/ SparseGPT 80.20 78.12 73.80 44.40 65.70 82.20 75.72 51.71 DSNoT w/ Wanda 79.82 77.99 73.09 44.80 63.18 81.80 77.06 51.37 OATS 80.03 78.75 73.64 45.20 66.06 81.13 76.94 52.99 40% SparseGPT 79.16 76.74 73.32 41.80 64.26 81.31 74.71 49.32 Wanda 78.73 75.90 72.22 44.40 63.18 80.46 72.31 49.15 DSNoT w/ SparseGPT 78.29 75.92 73.32 42.60 58.48 80.86 73.11 47.70 DSNoT w/ Wanda 78.51 75.52 73.24 43.80 61.73 80.70 72.01 47.70 OATS 79.71 77.18 74.19 43.80 67.51 82.39 74.92 49.74 50% SparseGPT 77.58 73.12 72.85 40.80 59.21 79.30 69.28 45.14 Wanda 77.53 69.34 70.24 40.00 61.73 76.57 66.96 43.77 DSNoT w/ SparseGPT 76.88 69.45 69.30 39.60 59.21 77.25 67.93 43.32 DSNoT w/ Wanda 77.09 68.57 69.77 38.60 57.76 76.27 67.34 43.43 OATS 77.75 73.17 71.74 41.00 64.98 79.66 72.35 45.05
Table 9: Task-Specific Zero-Shot Results
A.2 MMLU Subjects

### v2

@@ 3.3 Studies and Hyperparameter Exploration
The results are shown in Table 6 below:
@@
Table 6: Ablation results of OATS on Phi-3-Mini, at $40\%$ compression rate, with a rank ratio of $20\%$. Scaling the weights by the second moment of the input activations and pruning row-wise significantly improves the performance of OATS.
In addition to the ablations, we perform additional experiments to examine the impact of the rank ratio and the number of iterations on the performance of OATS. Figure 1 below shows the results.
0.10.20.30.40.50.60.740404040505050506060606070707070Rank RatioAccuracy (%)120406080100120140160Number of Iterations0-shotMMLU
@@
The experiments reveal that a rank ratio between $25\%$ to $30\%$ leads to the best performance, with degradation occurring at higher rank ratios. For the number of iterations, performance improves sharply in the first $20$ iterations, before leveling off and saturating at around $80$ iterations.
@@ 3.4 Hardware Speedup
We benchmark, using the DeepSparse engine by NeuralMagic (2021), the CPU throughput induced by OATS compared to models pruned with unstructured sparsity. We run end-to-end inference on a compressed Phi-3 Medium 15B model for a single batch of 2048 tokens on an Intel Xeon Gold 6148 CPU @ 2.40GHz with 32 cores. The achieved throughput and speedup (over a dense model) are shown in Table 7 below. By trading unstructured sparsity for structured sparsity through the low-rank terms, OATS achieves greater CPU speed-up compared to methods that rely solely on unstructured pruning. Notably, at $40\%$ compression, OATS is 1.37×1.37$\times$1.37 × faster than unstructured pruning.
Compression Method Throughput Speedup
0% Dense 4.03 1.00×1.$0$
@@
Table 7: Comparison of throughput (tokens/second) and speedup achieved through OATS and unstructured pruning methods relative to their dense counterparts.
@@
We compare the performance of state-of-the-art pruning algorithms, using a 2:4 structured sparsity pattern, with the performance of OATS, using a 2:8 structured sparsity pattern on the sparse term. OATS employs a sparser N:M pattern to compensate for its low-rank term that remains dense. We experiment with rank ratios of $\{0.25,0.3,0.35,0.4,0.45,0.5\}$. Unlike previous pruning methods, where N:M structured sparsity enforces a fixed compression rate of $\frac{N}{M}$, OATS allows for a flexible trade-off between compression and model performance by adjusting the rank ratio. Figure 2, below, illustrates the compression ratio against the 5-shot MMLU accuracy for various compression algorithms.
50545861646620202020303030304040404050505050Compression RateMMLU Avg. AccuracyPhi-3 Mini505458616466Compression RateLlama-3 8BOATSSparseGPTWandaDSNoT
@@
Despite having a sparser structured sparsity pattern of 2:8, OATS is able to recover the model performance through the presence of its low-rank term. Specifically, at a compression rate of $50\%$, OATS is able to outperform all prior state-of-the-art by $6.34\%$ on Phi-3 Mini. In the case of Llama-3 8B, OATS not only surpasses previous methods by $9.2\%$ at $50\%$ compression, but it also outperforms them by $2.86\%$ at an even higher compression rate of $54\%$.
@@ 4 Experiments on Vision Transformers
We run experiments on Google’s ViT-Base (Wu et al., 2020), an $86.6$M parameter model trained in a supervised manner on ImageNet-21k (Ridnik et al., 2021) and fine-tuned on ImageNet 2012 (Russakovsky et al., 2015), and DinoV2-Giant (Oquab et al., 2023), a $1.14$B parameter model that was trained through self-supervised learning.
We benchmark OATS against the same three pruning algorithms: SparseGPT, Wanda, and DSNoT, by evaluating top-1 accuracy on the validation set of ImageNet (Russakovsky et al., 2015). A subset of $2048$ images from the training set of ImageNet is used for calibration and is maintained consistent across all pruning experiments. All OATS experiments use a rank ratio of $\kappa{=}20\%$ and $N{=}80$ iterations. We exclude from compression the embedding and the classifier layers.
The results are shown in Table 8 below. Compared to LLMs, vision transformers show greater resilience to pruning, with DinoV2 experiencing only a $0.41\%$ drop in top-1 accuracy when compressed by $50\%$ using OATS.
Compression Method ViT-Base DinoV2-Giant
@@ 5 Visualizing and Interpreting the Decomposition
To develop a better understanding of how the sparse and low rank components individually contribute to the flow of information through the model, we compute and visualize the attention rollout (Abnar & Zuidema, 2020) of the compressed vision transformers when:
@@
Figure 3: A visualization of how the attention rollout is computed to isolate the contribution of the sparse terms versus low-rank terms given by the OATS algorithm.
Figure 4 depicts the attention rollout for various images in the Microsoft COCO dataset (Lin et al., 2014) passed to a ViT-B that was compressed by $50\%$, with a rank ratio of $20\%$.
@@
The rollout visualizations show that the sparse and low-rank terms capture distinct areas of the image, effectively segmenting it. A careful analysis reveals three distinct partitioning patterns. The first, which is also commonly exhibited in the classical setting (Candès et al., 2011), is when one component (commonly sparse) captures the subject(s), while the other component (commonly low-rank) captures the background. The second is when both components focus on different parts of the same subject, each capturing distinct features. The third behavior arises when the image contains multiple subjects, with each component isolating a different subject. While these patterns provide initial insights into how the components process visual information, further investigation is needed to fully understand the mechanisms driving these behaviors.
@@ 6 Related Works
OATS utilizes the same outlier scaling as the the one that is employed by Wanda (Sun et al., 2024b). In fact, Wanda can be seen as a special case of OATS when the rank ratio $\kappa{=}0$. Indeed, in such a case, according to Equation 2, the low-rank term would become the zero matrix and OATS would perform a single hard thresholding step that is equivalent to the pruning step described by Wanda: ${\bm{W}}_{\text{compressed}}=\textsc{HardThreshold}({\bm{W}}{\bm{D}},k){\bm{D}}^{-1}$.
@@
The emergence of sparse and low-rank structures in transformers has recently become an area of both theoretical and practical interest. On the theoretical front, Zhao et al. (2024c) showed that the logits of LLMs trained utilizing next token prediction converge to a low rank and sparse structure. On the practical front, Scatterbrain proposed by Chen et al. (2021a) shows that it is possible to approximate the entire attention mechanism with a single sparse and low rank decomposition. Pruning-wise, LoRAP by Li et al. (2024) performs structured pruning on the feed-forward linear layers and apply a low-rank decomposition to the attention matrices using a scaling technique similar to OATS.
Structured Pruning and Low-Rank Adaptation
Recent works, such as LoSparse (Li et al., 2023), LoRAPrune (Zhang et al., 2024a), and APT (Zhao et al., 2024a), propose variations of applying structured pruning on the weights while incorporating a low-rank adapter that is trained via gradient descent. These are markedly different than OATS, which does not employ any fine-tuning with low-rank adapters, nor does it perform structured pruning (but rather a sparse plus low-rank decomposition which can be thought of as a combination of structured and unstructured pruning).
@@
The search for Robust PCA algorithms has been a key area of interest since the inception of the problem. Examples of other approaches include applying a convex relaxation, where the sparsity and low-rank constraints are replaced by $\ell_{1}$ and nuclear norm surrogates (Zhou et al., 2010), or parameterizing the low-rank matrix as ${\bm{L}}={\bm{U}}{\bm{V}}^{\top}$, and applying gradient descent on ${\bm{U}}$ and ${\bm{V}}$ (Yi et al., 2016; Tong et al., 2021). While OATS utilizes the alternating thresholding approach for its simplicity, future work might want to investigate the use of other algorithms.
@@
An active area of research is understanding what pruning is pruning and how it impacts model performance. Paganini (2020) show that pruning has a disproportionate negative effect on underrepresented classes. In a similar vein, Yin et al. (2024a) showed that pruning LLMs can irreversibly harm model performance on tasks that are more challenging. We postulate that the low-rank term present in OATS might be able to mitigate the negative impacts of pruning. Indeed, Tables 2 and 5 show that the gap between OATS and prior methods is larger at higher compression, suggesting that the low-rank term plays a critical role in mitigating the loss in performance.
@@ 7 Conclusion
We have introduced OATS, an algorithm that without any re-training, compresses the model’s weight matrices through a sparse and low-rank decomposition. Taking inspiration from prior works on the emergence of outlier features, OATS first scales the weights by the second moment of their input embeddings prior to applying an alternating thresholding algorithm. A comprehensive evaluation shows that OATS is able to consistently outperform prior state-of-the-art on various performance metrics across multiple compression rates, models, and modalities, while also improving on CPU speed-up. Beyond just model compression, our visualizations on vision transformers indicate that models exhibit sparse and low-rank structures that capture different segments of the image. This work is the first to reveal the potential of sparse and low-rank decompositions for large-scale transformers, setting the stage for future innovations that can harness this structure to improve model efficiency, performance, and interpretability.
Acknowledgments
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). This research was supported in part by the Province of Ontario, the Government of Canada through CIFAR, and industry sponsors of the Vector Institute (www.vectorinstitute.ai/partnerships/current-partners/). This research was also enabled in part by support provided by Compute Ontario (https://www.computeontario.ca) and the Digital Research Alliance of Canada (https://alliancecan.ca).
@@ Appendix A Appendix
A.1 Additional Related Works
Sparse and Low-Rank Decomposition for Pruning
Yu et al. (2017) introduced a method for sparse and low-rank decomposition of CNNs, including AlexNet and GoogLeNet, by solving the following optimization problem:
$\min_{{\bm{S}},{\bm{L}}\in\mathbb{R}^{d_{out}\times d_{in}}}\|{\bm{Y}}-({\bm{S}}+{\bm{L}}){\bm{X}}\|_{2}^{2}\;\text{s.t.}\;\|{\bm{W}}-({\bm{S}}+{\bm{L}})\|_{F}^{2}\leq\gamma,\text{Rank}({\bm{L}})\leq r,\|{\bm{S}}\|_{0}\leq k$
where ${\bm{Y}}={\bm{W}}{\bm{X}}$. In contrast, OATS employs a different approach, solving:
$\min_{{\bm{S}},{\bm{L}}\in\mathbb{R}^{d_{out}\times d_{in}}}\|{\bm{W}}-{\bm{S}}-{\bm{L}}\|_{F}^{2}\;\text{ s.t. }\;\text{Rank}({\bm{L}})\leq r,\;\|{\bm{S}}\|_{0}\leq k.$
A key distinction between these methods lies in their objectives: the former directly minimizes reconstruction error, while OATS adopts a simpler formulation. One might question why not follow the approach of minimizing reconstruction error. As noted in DSNoT (Zhang et al., 2024b), pruning methods that prioritize minimizing reconstruction error can degrade model performance in large transformers, particularly in the presence of outlier features. Their findings highlight the importance of avoiding pruning weights within outlier channels. Since feature outliers are a phenomenon unique to large transformer models (Dettmers et al., 2022), this issue would not have been relevant to the work of Yu et al. (2017), which predates the transformer era.
Pruning Algorithms for Vision Transformers
There are a number of pruning approaches that have been specifically catered towards pruning vision transformers (Zhu et al., 2021; Chen et al., 2021b; Chavan et al., 2022; Yu et al., 2022a; b; Yu & Xiang, 2023). However, as much of the pruning literature developed on vision transformers involved models of much smaller scale than the large language models employed in this study, almost all of the prominent pruning algorithms require some form of training on the model parameters. As OATS was designed to require no training, OATS and the aforementioned pruning algorithms would not be comparable.
Low-Rank Adapters during Pre-Training
In Mozaffari et al. (2024), the authors propose SLOPE, a novel method for accelerating the pre-training phase of LLMs by incorporating N:M sparsity and adding low-rank components to the model weights to enhance model capacity. Similar to OATS, SLOPE leads to a sparse plus low-rank structure in the model’s weight matrices, however, the low-rank terms are introduced during the final phase of pre-training and are actively trained on the model loss function. In contrast, OATS is designed as a lightweight method to accelerate inference. OATS does not require any training or fine-tuning, but instead approximates pre-trained weight matrices by solving the Robust PCA problem.
Quantized Sparse Low-Rank Approximation
An independent and concurrent work with OATS proposes SLIM (Mozaffari & Dehnavi, 2024), a novel pipeline that combines pruning and quantization. To restore lost performance from compression, SLIM derives a low-rank term using singular-value thresholding and adopts a scaling technique akin to OATS. However, instead of the $L^{2}$ norm, SLIM utilizes the average absolute value across the batch and sequence dimensions. As a further deviation from OATS, SLIM is also not performing an alternating thresholding algorithm. Instead, they perform a single quantization and pruning step to initialize the quantized and sparse terms, followed by a single singular value thresholding step to establish the low-rank term.
A.2 Time Complexity and Wall-Clock Time for OATS
The time complexity for OATS is $\mathcal{O}(LN\alpha)$ where $L$ is the number of transformer blocks, $N$ is number of iterations, and
$\alpha=\max_{{\bm{W}}}d^{{\bm{W}}}_{out}\cdot d^{{\bm{W}}}_{in}\cdot r^{{\bm{W}}}$
where the max is taken over the weight matrices, ${\bm{W}}\in\mathbb{R}^{d^{{\bm{W}}}_{out}\times d^{{\bm{W}}}_{in}}$, in a transformer block and $r^{{\bm{W}}}$ is the rank of the low-rank term for that weight matrix. The value $\alpha$ represents the time complexity needed to perform the singular value thresholding in OATS.
Table 9 below reports the wall-clock time needed to perform a single iteration of the alternating threshold algorithm for a single transformer block for the different models that were compressed. All experiments utilized a single NVIDIA A40 with 48GB of GPU memory.
Phi-3 Llama-3
Mini ($\mathbf{3.8}$B) Medium ($\mathbf{14}$B) 8B 70B
8.85 26.02 17.10 152.80
Table 9: Wall-clock time (in seconds) needed to perform a single iteration of the alternating projection algorithm in OATS.
While OATS does require more wall-clock time than prior pruning algorithms, in practice, model compression would only need to be performed once before deployment. This trade-off is therefore worthwhile given the substantial performance improvements, particularly on more challenging tasks like MMLU (see Table 2). Furthermore, like prior pruning algorithms, compressing the layers within a single transformer block can be done in parallel. For example, the time needed per transformer block of Llama-3 70B can be reduced to 71.10 seconds by compressing in parallel across four NVIDIA A40 GPUs.
The total wall-clock time can also be reduced by lowering the number of OATS iterations. Presented in Table 10 is an exploratory experiment compressing Llama-3 70B by 50% with a rank ratio of 0.3 with only 20 iterations. Even with only a quarter of the iterations, OATS is still able to outperform all prior pruning algorithms across all performance metrics.
MMLU ()($\uparrow$)( ↑ ) Zero-shot ()($\uparrow$)( ↑ ) Perplexity ()($\downarrow$)( ↓ )
74.02 73.41 4.95
Table 10: Exploratory experiment measuring the performance of OATS on Llama-3 70B with only 20 iterations.
A.3 Using a Robust Scaling Matrix
To explore whether the scaling matrix ${\bm{D}}$ is truly related to the outlier information, we run the following two experiments:
  • Scaling by the square root of the features’ second moments, as is currently done in OATS.
  • Scaling by the median of the features’ absolute values (computed along batch and sequence dimensions):
    ${\bm{D}}_{robust}=\text{median}(|X|)$
The second experiment estimates the square root of the second moment of features in a manner that is robust (insensitive) to outliers akin to the Median Absolute Deviation estimator from the robust statistics literature (Huber, 1981). The results of the two experiments are presented in Table 11 below:
Scaling Matrix MMLU ()($\uparrow$)( ↑ ) Zero-shot ()($\uparrow$)( ↑ ) Perplexity ()($\downarrow$)( ↓ )
${\bm{D}}_{robust}$ 55.54 65.77 18.59
${\bm{D}}$ 59.99 68.41 15.18
Table 11: Results of OATS on Phi-3-Mini, at $50\%$ compression rate, with a rank ratio of $25\%$ using different scaling matrices.
The findings show that using the robust scaling method results in significantly worse performance. Hence, the scaling matrix ${\bm{D}}$ that is sensitive to the outlier features and captures their scale leads to better compression.
A.4 Switching the Order of Thresholding
OATS opts to perform the singular-value thresholding first followed by the hard thresholding similar to Zhou & Tao (2011). However, one might consider whether the alternative order could lead to faster convergence or a better approximation. Presented in Table 12 below is an extension of the ablation studies presented in Section 3.3, reporting the performance of OATS where the hard-thresholding is performed first:
First Thresholding Operation MMLU ()($\uparrow$)( ↑ ) Zero-shot ()($\uparrow$)( ↑ ) Perplexity ()($\downarrow$)( ↓ )
Hard-Thresholding 65.51 70.54 11.72
Singular Value Thresholding (OATS) 65.84 70.71 11.50
Table 12: Ablation results of switching of the order between the two thresholding operations. Experiments were run on Phi-3-Mini, at $40\%$ compression rate, with a rank ratio of $20\%$.
While the performance still remains competitive, across all performance metrics, the switched order falls short of matching the original order presented in Algorithm 1.
A.5 Magnitude-Based Pruning for the Sparse Component
Another question that we explored is whether it is sufficient to capture the outlier information entirely in the low-rank term and determine the sparse term through a hard-thresholding that does not depend on the scaling:
${\bm{S}}=\textsc{HardThreshold}(({\bm{W}}{\bm{D}}-{\bm{L}}){\bm{D}}^{-1},k).$
Presented in Table 13 below are the results:
Outlier Scaling MMLU ()($\uparrow$)( ↑ ) Zero-shot ()($\uparrow$)( ↑ ) Perplexity ()($\downarrow$)( ↓ )
Low-Rank Term Only 65.22 71.01 12.49
Both Terms (OATS) 65.84 70.71 11.50
Table 13: Ablation results of OATS on Phi-3-Mini, at $40\%$ compression rate, with a rank ratio of $20\%$ testing whether the outlier information can be entirely captured by the low-rank term.
A.6 Additional Hyperparameter Tests for OATS
Presented in Table 14 below includes more hyperparameters that we experimented with for the Phi-3 Mini and Llama-3 8B models.
Model Compression Rank Ratio MMLU ($\uparrow$) Zero-Shot ($\uparrow$) Perplexity ($\downarrow$)
Phi-3 Mini 30% 0.1 68.70 71.65 10.24
0.2 68.02 71.81 10.21
0.3 69.28 72.07 10.28
40% 0.1 65.75 69.94 11.57
0.2 65.84 70.71 11.50
0.3 66.81 70.54 11.60
50% 0.1 57.96 67.37 15.48
0.2 59.12 68.02 15.13
0.3 58.68 68.63 15.47
Llama-3 8B 30% 0.1 63.62 68.99 9.35
0.2 63.09 69.54 9.09
40% 0.1 61.44 68.23 9.23
0.2 61.97 68.43 9.09
50% 0.1 56.46 65.33 10.85
0.2 56.07 65.51 10.70
Table 14: Further experiments testing different hyperparameter configurations for OATS on the Phi-3 Mini and Llama-3 8B models.
A.7 Performance Gap Between OATS and Wanda
To better understand the increase in performance induced by the addition of the low-rank term in OATS, we have compiled in Table 15 below the performance gaps between OATS and Wanda.
Model Compression MMLU ($\uparrow$) Zero-Shot ($\uparrow$) Perplexity ($\downarrow$)
Phi-3 Mini 30% +1.21% +0.82% -0.44
40% +1.60% +1.24% -1.06
50% +5.42% +3.38% -2.05
Phi-3 Medium 30% +0.97% -0.01% -0.43
40% +1.65% +1.45% -0.79
50% +2.52% +2.43% -1.07
Llama-3 8B 30% +1.55% +0.71% +0.20
40% +2.13% +1.64% -0.50
50% +6.63% +2.44% -1.49
Llama-3 70B 30% -0.68% +0.05% -0.17
40% +0.73% +0.78% -0.40
50% +2.74% +0.45% -0.60
Table 15: The impact of including a low-rank term in OATS compared to Wanda.
A.8 Qwen 2.5 Experiments
Presented in Table 16 below are additional experiments benchmarking OATS against prior pruning algorithms on the Qwen 2.5 3B Instruct model (Qwen Team, 2024). All OATS experiments utilize a rank ratio of $0.2$ and $80$ iterations.
Compression Method MMLU ($\uparrow$) Zero-Shot ($\uparrow$) Perplexity ($\downarrow$)
$0\%$ Dense 65.99 68.49 11.02
30% SparseGPT 65.65 67.91 11.55
Wanda 65.46 68.08 11.66
DSNoT 65.65 68.21 11.67
OATS 65.36 68.74 11.45
40% SparseGPT 63.04 67.64 12.56
Wanda 61.88 67.14 12.89
DSNoT 62.26 67.42 12.91
OATS 64.30 68.76 12.31
50% SparseGPT 57.43 64.36 14.92
Wanda 55.39 64.10 16.27
DSNoT 55.78 64.77 16.43
OATS 58.78 65.74 14.91
Table 16: Benchmarks for OATS on the Qwen 2.5 3B Instruct model.
A.9 MMLU Subjects